OpenStudy (anonymous):

let's say i have (x^2 . y^3).2.5 how do i get rid of the brackets?

OpenStudy (aum):

Since there is just multiplication inside the bracket, it can be removed without affecting anything. \( (x^2 . y^3).2.5 = x^2 . y^3 * 2.5 = 10x^2y^3\)

OpenStudy (aum):

I am assuming it is 2 * 5 and not 2.5

OpenStudy (anonymous):

i'm sorry for my bad notation, it was 2,5 but thnx!!

OpenStudy (aum):

2.5x^2y^3